# Blind Inverse Game Theory: Jointly Decoding Rewards and Rationality in Entropy-Regularized Competitive Games

### v1

@@ C.3 Proof of Theorem 2 (Finite Sample Bounds)
$\|\hat{\theta}-\theta^{*}\|_{2}$ $=\left\|\frac{\hat{\theta}_{LS}}{\|\hat{\theta}_{LS}\|_{2}}-\frac{\theta^{\prime*}}{\|\theta^{\prime*}\|_{2}}\right\|_{2}$ (70)
$\leq\frac{4\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}}{\|\theta^{\prime*}\|_{2}}=\frac{4C_{LS}\epsilon_{N}}{1/\tau^{*}}=4\tau^{*}C_{LS}\epsilon_{N}.$ (71)
@@
$|\hat{\tau}-\tau^{*}|$ $=\left|\frac{1}{\|\hat{\theta}_{LS}\|_{2}}-\frac{1}{\|\theta^{\prime*}\|_{2}}\right|=\frac{|\|\theta^{\prime*}\|_{2}-\|\hat{\theta}_{LS}\|_{2}|}{\|\hat{\theta}_{LS}\|_{2}\|\theta^{\prime*}\|_{2}}.$ (72)
@@
$|\hat{\tau}-\tau^{*}|$ $\leq\frac{C_{LS}\epsilon_{N}}{(\|\theta^{\prime*}\|_{2}/2)\|\theta^{\prime*}\|_{2}}=\frac{2C_{LS}\epsilon_{N}}{\|\theta^{\prime*}\|_{2}^{2}}$ (73)
$=2(\tau^{*})^{2}C_{LS}\epsilon_{N}.$ (74)
@@
C.4 Proof of Proposition 1 (Partial Identification)
@@
We now substitute the bounds established in the supporting lemmas. Let $\epsilon_{N}$ be the convergence rate of the policy estimation (Lemma 1). From Lemma 2, we know that with probability at least $1-\delta$ (assuming $N$ is sufficiently large such that the conditions of the lemma hold):
@@
By Assumption 4, we have $\tau^{*}\leq\tau_{max}$. Therefore, the residual is uniformly bounded over $\Theta^{*}$:
@@
Appendix D Proofs for Markov Games
This section provides the detailed proofs for the theoretical results presented in Section 5 concerning Blind-IGT in Markov Games. We assume the dynamics $P$ are known, as stated in Theorem 3. We assume Assumption 3 holds uniformly for the QRE policies at every state $s\in\mathcal{S}$, i.e., $\min_{s,a,b}\{\mu^{*}(a|s),\nu^{*}(b|s)\}\geq\xi$.
D.1 Proof of Proposition 2 (Identifiability in MGs)
@@
The mathematical structure of this aggregated bilinear system is identical to the matrix game formulation analyzed in Theorem 1. Therefore, the necessary and sufficient conditions for unique identification directly translate. The pair $(\theta^{*},\tau^{*})$ is uniquely identifiable if and only if:
@@
The proof follows exactly the steps outlined in Appendix C.1, applied to the aggregated matrices $X^{*}$ and $y^{*}$. ∎
D.2 Proof of Theorem 3 (Sample Complexity for MGs)
The proof builds upon the matrix game analysis (Appendix C.3) and incorporates the complexities of dynamic programming, specifically the error propagation during reward recovery.
@@ D.2 Proof of Theorem 3 (Sample Complexity for MGs)
Uniform Policy Estimation Error. We first establish the uniform convergence of the QRE policy estimates $(\hat{\mu}(s),\hat{\nu}(s))$. We observe samples drawn from the stationary distribution $d^{*}(s)$ induced by the QRE. Under Assumption 6 ($d^{*}(s)\geq C_{\min}>0$), the expected number of visits to state $s$ is approximately $K\cdot d^{*}(s)\geq KC_{\min}$.
We utilize concentration results for empirical estimates in Markov chains (or assume access to a generative model allowing sampling from $d^{*}$). This ensures that the empirical estimators converge uniformly across all states $s\in\mathcal{S}$. By adapting the results from Lemma 1 and applying a union bound over all states, we obtain that with probability at least $1-\delta$:
$\sup_{s}(\|\hat{\mu}(s)-\mu^{*}(s)\|_{1}+\|\hat{\nu}(s)-\nu^{*}(s)\|_{1})=\mathcal{O}\left(\sqrt{\frac{(m+n)\log(|\mathcal{S}|mn/\delta)}{KC_{\min}}}\right).$ (75)
@@
$\|\hat{X}-X^{*}\|_{op}$ $\leq\|\hat{X}-X^{*}\|_{F}=\sqrt{\sum_{s}\|\hat{X}(s)-X^{*}(s)\|_{F}^{2}}.$ (76)
Using the analysis from Lemma 2, the local perturbation $\|\hat{X}(s)-X^{*}(s)\|_{F}$ is bounded by $C_{X}(s)\epsilon_{K}(s)$, where $\epsilon_{K}(s)$ is the local policy error at state $s$. Since we established uniform convergence $\epsilon_{K}$, and $C_{X}$ depends on $L,m,n$ (which are uniform across states), we have:
$\|\hat{X}-X^{*}\|_{op}\leq\sqrt{|\mathcal{S}|C_{X}^{2}\epsilon_{K}^{2}}=\sqrt{|\mathcal{S}|}C_{X}\epsilon_{K}.$ (77)
@@
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (78)
Applying the normalization stability (Lemma 3), we obtain the bounds for the parameters:
$\|\hat{\theta}-\theta^{*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}),\quad|\hat{\tau}-\tau^{*}|=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (79)
@@
$\hat{r}(s,a,b)=\hat{Q}(s,a,b)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[\hat{V}(s^{\prime})].$ (80)
@@
$\|\hat{Q}-Q^{*}\|_{\infty}=\sup_{s,a,b}|\langle\phi(s,a,b),\hat{\theta}-\theta^{*}\rangle|\leq L\|\hat{\theta}-\theta^{*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (82)
@@
We analyze the Lipschitz continuity of the function $F$. Provided the policies are bounded away from zero (by $\xi>0$, Assumption 3), $F$ is Lipschitz continuous with respect to its arguments.
@@
$\|\hat{V}-V^{*}\|_{\infty}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (84)
@@
Appendix E Experimental Details and Additional Results
E.1 Experimental Setup Details
@@
Matrix Games Generation. For the matrix game experiments (Sections 6.2, 6.3), we set the dimensions $m=10,n=10$ and the feature dimension $d=5$. The features $\phi(a,b)\in\mathbb{R}^{d}$ were generated by sampling each element independently from a standard normal distribution $\mathcal{N}(0,1)$. The true reward parameter $\theta^{*}\in\mathbb{R}^{d}$ was also sampled from $\mathcal{N}(0,1)$ and then normalized such that $\|\theta^{*}\|_{2}=1$ (Normalization constant $C=1$), unless otherwise specified (e.g., Section 4.5). The true temperature was fixed at $\tau^{*}=2.0$. The payoff matrix $Q^{*}$ was computed as $Q^{*}(a,b)=\langle\phi(a,b),\theta^{*}\rangle$. The QRE $(\mu^{*},\nu^{*})$ was computed by solving the fixed-point equations (Eq. 2, 3) using iterative updates until convergence (tolerance $10^{-9}$), falling back to a numerical root-finding solver if iterations failed. Samples were drawn directly from the computed QRE.
Markov Games Generation. For the Markov game experiments (Sections 6.2, 6.4), we set $|\mathcal{S}|=8,m=5,n=5,d=6$, and $\gamma=0.9$. We adopted the Linear Q-function parameterization (Assumption 5). Features $\phi(s,a,b)\in\mathbb{R}^{d}$ were sampled from $\mathcal{N}(0,1)$. The true parameter $\theta^{*}$ was sampled from $\mathcal{N}(0,1)$ and normalized to $\|\theta^{*}\|_{2}=1$ (Normalization constant $R=1$). The true temperature was $\tau^{*}=1.5$. The QRE Q-function $Q^{*}$ was defined as $Q^{*}(s,a,b)=\langle\phi(s,a,b),\theta^{*}\rangle$. The QRE policies $(\mu^{*},\nu^{*})$ were computed from $Q^{*}$ using the logit response equations (using the same solver approach as Matrix Games). The V-function $V^{*}$ was computed using Eq. 6.
The transition dynamics $P(s^{\prime}|s,a,b)$ were generated randomly by sampling from a Dirichlet distribution $\text{Dir}(1)$ for each $(s,a,b)$. The true reward function $r^{*}$ was then computed by inverting the Bellman equation (Eq. 5): $r^{*}(s,a,b)=Q^{*}(s,a,b)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[V^{*}(s^{\prime})]$.
@@
E.2 Robustness to Feature Misspecification
A critical assumption in IGT/IRL is that the feature map $\phi$ (Assumption 1) perfectly captures the relevant factors influencing the agents’ utilities. We analyze the robustness of Blind-IGT when this assumption is violated, specifically when the feature map used for estimation is incomplete.
E.2.1 Setup

### v2

@@ D.2 Supporting Lemmas for Finite Sample Analysis
$\|\hat{\theta}-\theta^{*}\|_{2}$ $=\left\|\frac{\hat{\theta}_{LS}}{\|\hat{\theta}_{LS}\|_{2}}-\frac{\theta^{\prime*}}{\|\theta^{\prime*}\|_{2}}\right\|_{2}$ (D.41)
$\leq\frac{4\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}}{\|\theta^{\prime*}\|_{2}}=\frac{4C_{LS}\epsilon_{N}}{1/\tau^{*}}=4\tau^{*}C_{LS}\epsilon_{N}.$ (D.42)
@@
$|\hat{\tau}-\tau^{*}|$ $=\left|\frac{1}{\|\hat{\theta}_{LS}\|_{2}}-\frac{1}{\|\theta^{\prime*}\|_{2}}\right|=\frac{|\|\theta^{\prime*}\|_{2}-\|\hat{\theta}_{LS}\|_{2}|}{\|\hat{\theta}_{LS}\|_{2}\|\theta^{\prime*}\|_{2}}.$ (D.43)
@@
$|\hat{\tau}-\tau^{*}|$ $\leq\frac{C_{LS}\epsilon_{N}}{(\|\theta^{\prime*}\|_{2}/2)\|\theta^{\prime*}\|_{2}}=\frac{2C_{LS}\epsilon_{N}}{\|\theta^{\prime*}\|_{2}^{2}}$ (D.44)
$=2(\tau^{*})^{2}C_{LS}\epsilon_{N}.$ (D.45)
@@
D.4 Proof of Proposition 4.8 (Partial Identification)
@@
We now substitute the bounds established in the supporting lemmas. Let $\epsilon_{N}$ be the convergence rate of the policy estimation (Lemma D.1). From Lemma D.2, we know that with probability at least $1-\delta$ (assuming $N$ is sufficiently large such that the conditions of the lemma hold):
@@
By Assumption 4.7, we have $\tau^{*}\leq\tau_{max}$. Therefore, the residual is uniformly bounded over $\Theta^{*}$:
@@
Appendix E Proofs for Markov Games
This section provides the detailed proofs for the theoretical results presented in Section 5 concerning Blind-IGT in Markov Games. We assume the dynamics $P$ are known, as stated in Theorem 5.4. We assume Assumption 4.5 holds uniformly for the QRE policies at every state $s\in\mathcal{S}$, i.e., $\min_{s,a,b}\{\mu^{*}(a|s),\nu^{*}(b|s)\}\geq\xi$.
E.1 Proof of Proposition 5.2 (Identifiability in MGs)
@@ Appendix E Proofs for Markov Games
The mathematical structure of this aggregated bilinear system is identical to the matrix game formulation analyzed in Theorem 4.4. Therefore, the necessary and sufficient conditions for unique identification directly translate. The pair $(\theta^{*},\tau^{*})$ is uniquely identifiable if and only if:
@@
The proof follows exactly the steps outlined in Appendix D.1, applied to the aggregated matrices $X^{*}$ and $y^{*}$. ∎
E.2 Proof of Theorem 5.4 (Sample Complexity for MGs)
The proof builds upon the matrix game analysis (Appendix D.3) and incorporates the complexities of dynamic programming, specifically the error propagation during reward recovery.
@@
Uniform Policy Estimation Error. We first establish the uniform convergence of the QRE policy estimates $(\hat{\mu}(s),\hat{\nu}(s))$. We observe samples drawn from the stationary distribution $d^{*}(s)$ induced by the QRE. Under Assumption 5.3 ($d^{*}(s)\geq C_{\min}>0$), the expected number of visits to state $s$ is approximately $K\cdot d^{*}(s)\geq KC_{\min}$.
We utilize concentration results for empirical estimates in Markov chains (or assume access to a generative model allowing sampling from $d^{*}$). This ensures that the empirical estimators converge uniformly across all states $s\in\mathcal{S}$. By adapting the results from Lemma D.1 and applying a union bound over all states, we obtain that with probability at least $1-\delta$:
$\sup_{s}(\|\hat{\mu}(s)-\mu^{*}(s)\|_{1}+\|\hat{\nu}(s)-\nu^{*}(s)\|_{1})=\mathcal{O}\left(\sqrt{\frac{(m+n)\log(|\mathcal{S}|mn/\delta)}{KC_{\min}}}\right).$ (E.1)
@@
$\|\hat{X}-X^{*}\|_{op}$ $\leq\|\hat{X}-X^{*}\|_{F}=\sqrt{\sum_{s}\|\hat{X}(s)-X^{*}(s)\|_{F}^{2}}.$ (E.2)
Using the analysis from Lemma D.2, the local perturbation $\|\hat{X}(s)-X^{*}(s)\|_{F}$ is bounded by $C_{X}(s)\epsilon_{K}(s)$, where $\epsilon_{K}(s)$ is the local policy error at state $s$. Since we established uniform convergence $\epsilon_{K}$, and $C_{X}$ depends on $L,m,n$ (which are uniform across states), we have:
$\|\hat{X}-X^{*}\|_{op}\leq\sqrt{|\mathcal{S}|C_{X}^{2}\epsilon_{K}^{2}}=\sqrt{|\mathcal{S}|}C_{X}\epsilon_{K}.$ (E.3)
@@
$\|\hat{\theta}_{LS}-\theta^{\prime*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (E.4)
Applying the normalization stability (Lemma D.3), we obtain the bounds for the parameters:
$\|\hat{\theta}-\theta^{*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}),\quad|\hat{\tau}-\tau^{*}|=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (E.5)
@@
$\hat{r}(s,a,b)=\hat{Q}(s,a,b)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[\hat{V}(s^{\prime})].$ (E.6)
@@
$\|\hat{Q}-Q^{*}\|_{\infty}=\sup_{s,a,b}|\langle\phi(s,a,b),\hat{\theta}-\theta^{*}\rangle|\leq L\|\hat{\theta}-\theta^{*}\|_{2}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (E.8)
@@
We analyze the Lipschitz continuity of the function $F$. Provided the policies are bounded away from zero (by $\xi>0$, Assumption 4.5), $F$ is Lipschitz continuous with respect to its arguments.
@@
$\|\hat{V}-V^{*}\|_{\infty}=\mathcal{O}(\sqrt{|\mathcal{S}|}\epsilon_{K}).$ (E.10)
@@
Appendix F Experimental Details and Additional Results
F.1 Experimental Setup Details
@@
Matrix Games Generation. For the matrix game experiments (Sections 6.2, 6.3), we set the dimensions $m=10,n=10$ and the feature dimension $d=5$. The features $\phi(a,b)\in\mathbb{R}^{d}$ were generated by sampling each element independently from a standard normal distribution $\mathcal{N}(0,1)$. The true reward parameter $\theta^{*}\in\mathbb{R}^{d}$ was also sampled from $\mathcal{N}(0,1)$ and then normalized such that $\|\theta^{*}\|_{2}=1$ (Normalization constant $C=1$), unless otherwise specified (e.g., Section 4.5). The true temperature was fixed at $\tau^{*}=2.0$. The payoff matrix $Q^{*}$ was computed as $Q^{*}(a,b)=\langle\phi(a,b),\theta^{*}\rangle$.
The QRE $(\mu^{*},\nu^{*})$ was computed by solving the fixed-point equations (Eq. 3.2, 3.3) using iterative updates until convergence (tolerance $10^{-9}$), falling back to a numerical root-finding solver if iterations failed. Samples were drawn directly from the computed QRE.
Markov Games Generation. For the Markov game experiments (Sections 6.2, 6.4), we set $|\mathcal{S}|=8,m=5,n=5,d=6$, and $\gamma=0.9$. We adopted the Linear Q-function parameterization (Assumption 5.1). Features $\phi(s,a,b)\in\mathbb{R}^{d}$ were sampled from $\mathcal{N}(0,1)$. The true parameter $\theta^{*}$ was sampled from $\mathcal{N}(0,1)$ and normalized to $\|\theta^{*}\|_{2}=1$ (Normalization constant $R=1$). The true temperature was $\tau^{*}=1.5$. The QRE Q-function $Q^{*}$ was defined as $Q^{*}(s,a,b)=\langle\phi(s,a,b),\theta^{*}\rangle$. The QRE policies $(\mu^{*},\nu^{*})$ were computed from $Q^{*}$ using the logit response equations (using the same solver approach as Matrix Games). The V-function $V^{*}$ was computed using Eq. 3.6.
The transition dynamics $P(s^{\prime}|s,a,b)$ were generated randomly by sampling from a Dirichlet distribution $\text{Dir}(1)$ for each $(s,a,b)$. The true reward function $r^{*}$ was then computed by inverting the Bellman equation (Eq. 3.5): $r^{*}(s,a,b)=Q^{*}(s,a,b)-\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a,b)}[V^{*}(s^{\prime})]$.
@@
F.2 Robustness to Feature Misspecification
A critical assumption in IGT/IRL is that the feature map $\phi$ (Assumption 4.1) perfectly captures the relevant factors influencing the agents’ utilities. We analyze the robustness of Blind-IGT when this assumption is violated, specifically when the feature map used for estimation is incomplete.
F.2.1 Setup